A particle of mass 10 gm is describing shm along a straight line with the period of 2sec and amplitude of 10 cm its kinetic energy when it is at 5 cm from its equillibrium position

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Pallavi

let equation of SHM is x=Xsinwt

wher w=2∏/T         and T= 2sec

X=10 cm

w=√k/m

where k is spring constant

k=w2m

  =(2∏/T)2m

   =∏2m

K.E at x=5 cm  +P.E at x=5  =P.E at x=10cm(extreme position)

  K.E at x=5 cm = 1/2 k 102  -  1/2 k52

                      =1/2 k(100 -25)

                      =1/2 ∏2m *75

      now solve


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed

solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin